Operation Summary for January 6, 2010 - January 8, 2010 |
Wednesday Jan. 6 The day shift began with Operations monitoring the Tevatron (TeV) store 7485 (85.72E30) and the Recycler stash (306.8E10), with the Antiproton Source stacking (11.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7485 at 1:30 PM. During the evening shift, Operations put the TeV into HEP shot setup and established store 7486 at 4:34 PM with an average initial luminosity (AIL) of 301.6E30. Thursday Jan. 7 The midnight shift began with Operations monitoring store 7486 (108.94E30) and the stash (267.5E10), with Pbar stacking (13.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7486 at 6:48 AM. Operations put the TeV into shot setup, but a problem with a RF station (TRF6) modulator halted the setup. During the day shift, Operations reported that shot setup was delayed due to a TRF6 crate power supply failure, not a modulator. After the supply was repaired, Operations established store 7487 at 10:04 AM with an AIL of 242.623E30. During the evening shift, Operations made regular transfers of antiprotons from Pbar to the Recycler and sent beam to Pbar, NuMI, and MiniBooNE as available. Friday Jan. 8 The midnight shift began with Operations monitoring store 7487 (45.25E30) and the stash (72.2E10), with Pbar stacking (276.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7487 at 7:03 AM and then prepared the CDF and D0 collision halls for accesses. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Pbar Debuncher cooling studies. NuMI reported breaking a best day record for protons on target with 13.42E17 and that last year they passed 10E20. |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |